Find the value of “p” from the polynomial x2 + 3x + p, if one of the zeroes of the polynomial is 2.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

As 2 is the zero of the polynomial.

We know that ifα is a zero of the polynomial p(x),
then p(α) = 0

Substituting x = 2 in x^2+ 3x + p,

⇒ 2^2+ 3(2) + p = 0

⇒ 4 + 6 + p = 0

⇒ 10 + p = 0

⇒ p = -10

Thanks